# EPO: Entropy-regularized Policy Optimization for LLM Agents Reinforcement Learning

### v1

@@ 4.1 Entropy Regularization
To address the exploration-exploitation cascade failure, we first adapt entropy regularization to capture the temporal structure of multi-turn interactions. Unlike traditional RL where entropy is computed per-step, we recognize that in multi-turn environments, early trajectory decisions compound across subsequent turns. Therefore, we compute entropy across all turns within each trajectory and average over the batch of trajectories. The entropy-regularized policy loss is formulated as $L^{ER}(\theta)=L^{MT}(\theta)-\lambda L^{H}(\theta)$, where $\lambda$ is the entropy coefficient, and the entropy loss is averaged over the batch of trajectories:
@@ 4.2 Entropy Smoothing Regularizer
To break the cascade failure’s two-phase pattern, excessive early exploration followed by late-stage uncertainty propagation, we introduce an entropy smoothing mechanism that prevents the dangerous oscillations observed in sparse reward settings. We maintain an entropy history window $\mathcal{W}_{k}=\{\bar{H}_{0},\ldots,\bar{H}_{m},\ldots,\bar{H}_{k-1}\}$ for RL step $k$, storing the average entropy $\bar{H}_{m}$ across all trajectories at the token level for each previous RL step $m$. The historical entropy reference is computed as the mean $\bar{H}^{\mathcal{W}_{k}}=\frac{1}{k}\sum_{m=0}^{k-1}\bar{H}_{m}$. This historical anchoring prevents the uncontrolled entropy growth that characterizes the early exploration phase of cascade failure. We apply a token-wise penalty based on acceptable entropy ranges relative to this historical average:
$\mathcal{P}_{n,t,i}=\begin{cases}0,&\text{if }\kappa_{l}\bar{H}^{\mathcal{W}_{k}}\leq\mathcal{H}_{n,t,i}\leq\kappa_{r}\bar{H}^{\mathcal{W}_{k}},\\ \alpha,&\text{otherwise.}\end{cases}$ (5)
The boundary coefficients $\kappa_{l}$ and $\kappa_{r}$ define the acceptable range, while $\alpha$ provides the penalty weight for tokens with entropy outside the desired range. By bounding entropy within historical averages, we prevent both the blind exploration of early stages and the chaotic uncertainty propagation of late stages. Aggregating these penalties across all tokens, turns, and trajectories yields the smoothing loss:
$L^{smooth}(\theta)=\frac{1}{B}\sum_{n=1}^{B}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{|\tau_{n,t}|}\sum_{i=1}^{|\tau_{n,t}|}\mathcal{P}_{n,t,i}$ (6)
@@
The complete entropy-smoothed policy optimization loss is then defined as $L^{EPO}(\theta)=L^{MT}(\theta)-\lambda[L^{H}(\theta)-\beta_{k}L^{smooth}(\theta)]$, where the dynamic coefficient $\beta_{k}$ follows an exponential schedule that directly counteracts the cascade failure’s progression:
$\beta_{k}=\begin{cases}\beta_{start}+(1-\beta_{start})\left(1-e^{-\lambda_{d}\frac{k}{k_{mid}}}\right),&\text{if }k\leq k_{mid}\\ 1+(\beta_{end}-1)\left(1-e^{-\lambda_{d}\frac{k-k_{mid}}{K-k_{mid}}}\right),&\text{if }k>k_{mid}\end{cases}$ (7)
This adaptive schedule specifically addresses each phase of the cascade failure: it begins with conservative exploration to prevent early-stage behavioral lock-in, transitions through balanced exploration-exploitation at mid-training ($k_{mid}=\lfloor K/2\rfloor$), and increases smoothing strength in later phases to prevent uncertainty propagation and ensure stable convergence. The parameters $\beta_{start}$, $\beta_{end}$, $\lambda_{d}$, and $K$ control the transition dynamics and training duration.
Algorithm 1 presents the full optimization procedure, incorporating these components.
Algorithm 1 Entropy-smoothed Policy Optimization (EPO)
@@
1:Initial policy parameters $\theta$, entropy coefficient $\lambda$, smoothing penalty $\alpha$, entropy bounds $\kappa_{l},\kappa_{r}$, and schedule parameters for $\beta_{k}$.
2:Initialize policy parameters $\theta_{0}$.
3:Initialize historical entropy window $\mathcal{W}_{0}=\emptyset$.
4:for each training step $k=0,1,2,\ldots,K-1$ do
5:  Collect a batch of $B$ trajectories $\mathcal{D}_{k}=\{\tau_{j}\}_{j=1}^{B}$ using the old policy $\pi_{\theta_{\text{old}}}$.
6:  Compute advantages $\hat{A}_{t}$ for all timesteps in $\mathcal{D}_{k}$.
7:  Compute the multi-turn policy loss $L^{MT}(\theta)$ using  Equation 3.
8:  $\triangleright$ — Start of EPO specific computations —
9:  Compute the trajectory-aware entropy loss $L^{H}(\theta)$ over the batch using  Equation 4.
10:  if $k>0$ then
11:   Compute the historical entropy mean $\bar{H}^{\mathcal{W}_{k}}$ from the window $\mathcal{W}_{k}$.
12:   Compute the smoothing loss $L^{\text{smooth}}(\theta)$ based on $\bar{H}^{\mathcal{W}_{k}}$ using  Equation 6.
13:  else
14:   $L^{\text{smooth}}(\theta)\leftarrow 0$.
15:  end if
16:  Update the dynamic coefficient $\beta_{k}$ according to the schedule in Equation 7.
17:  Combine losses to form the final EPO objective:
18:  $L^{EPO}(\theta)=L^{MT}(\theta)-\lambda[L^{H}(\theta)-\beta_{k}L^{\text{smooth}}(\theta)]$.
19:  Update the policy parameters $\theta$ by optimizing $L^{EPO}(\theta)$.
20:  Update the historical entropy window $\mathcal{W}_{k+1}\leftarrow\mathcal{W}_{k}\cup\{\bar{H}_{\text{current batch}}\}$.
21:end for
22:Optimized policy parameters $\theta_{K}$.
5 Experiments
5.1 Experiments Setup
Benchmark. We evaluate on two challenging benchmarks that require different reasoning capabilities, ScienceWorld (Wang et al., 2022) and ALFWorld (Shridhar et al., 2021). ScienceWorld focuses on text-based scientific experimentation, demanding systematic hypothesis testing and structured exploration. ALFWorld is an embodied environment containing 4,639 household task instances across six categories, requiring multi-step decision-making and spatial reasoning. To improve the generalizability of our approach across multiple scenarios, we finetune the foundation model directly on the environment using RL, rather than employing trajectory finetuning for initialization.
Baselines. We conduct comprehensive comparisons across multiple paradigms: (1) Prompting-based approaches such as ReAct (Yao et al., 2023) that utilize in-context learning without parameter optimization; (2) Trajectory-based and platform methods including supervised fine-tuning (SFT) through expert trajectory imitation and AgentGym (Xi et al., 2024) which provides a unified framework with behavioral cloning and self-evolution mechanisms; (3) General reinforcement learning approaches encompassing standard on-policy methods (PPO (Schulman et al., 2017), GRPO (Shao et al., 2024)); (4) Agent RL approaches including recent methods specifically designed for agent training (GiGPO (Feng et al., 2025), RLVMR (Zhang et al., 2025)). Our proposed EPO methodology is architected as a general enhancement framework that can be seamlessly integrated with existing RL paradigms, as exemplified through our PPO+EPO and GRPO+EPO implementations.
Implementation Details. To optimize performance for each benchmark’s complexity, we employ Qwen2.5-3B-Instruct for ALFWorld and the larger Qwen2.5-7B-Instruct for ScienceWorld’s more complex scientific reasoning tasks. Constrained by computational resources, we adopt a single foundation model per task whose size is sufficient to ensure proper convergence, as smaller models consistently fail to converge. We conduct our own implementations and experiments for our proposed method, PPO, and GRPO baselines across three random seeds to ensure statistical reliability. Results for other baseline methods (ReAct, AgentGym, SFT, GiGPO, RLVMR) are sourced from the RLVMR’s (Zhang et al., 2025) paper. To account for their different convergence characteristics, we trained the model for 120 RL steps on ScienceWorld and 150 steps on ALFWorld.
5.2 Performance Comparison
@@
Table 1: For PPO and GRPO baselines with our EPO method: better performance indicates improvement over baseline, worse performance indicates degradation. Highlighted values represent the best performance among other baseline methods. $\Delta$ shows the relative improvement (%) when applying our method. Results for other baseline methods (ReAct, AgentGym, SFT, GiGPO, RLVMR) are sourced from the RLVMR (Zhang et al., 2025) paper. We ran our own implementations of PPO, GRPO, and EPO, tuning hyperparameters across multiple trials to obtain stable results.
@@
ReAct GPT-4o 45.4 - 49.2 - GPT-4o 57.3 - 66.0 -
ReAct DeepSeek-R1 22.2 - 31.4 - DeepSeek-R1 68.8 - 70.2 -
ReAct Qwen2.5-7B 7.8 - 11.3 - Qwen2.5-7B 23.1 - 28.5 -
AgentGym LLaMa2-7B 46.9 - 33.6 - LLaMa2-7B 76.6 - 63.3 -
@@
GiGPO Qwen2.5-7B 53.4 - 35.2 - Qwen2.5-7B 89.5 - 90.2 -
RLVMR Qwen2.5-7B 67.2 - 43.0 - Qwen2.5-7B 91.4 - 91.8 -
PPO Qwen2.5-7B 64.6 38.4 58.3 39.1 Qwen2.5-3B 95.8 72.3 87.5 70.9
@@
GRPO Qwen2.5-7B 93.8 81.6 91.7 80.9 Qwen2.5-3B 87.5 63.3 83.3 63.5
@@
Figure 2: Training dynamics and generalization performance analysis. We present the evolution of training rewards and validation success rates across both in-distribution (IID) and out-of-distribution (OOD) evaluation settings. (a-c) ScienceWorld experimental results contrasting PPO and PPO+EPO performance across training reward accumulation, IID validation, and OOD validation metrics. (d-f) ALFWorld experimental results contrasting GRPO and GRPO+EPO under identical evaluation criteria. Our EPO enhancement exhibits significantly improved training stability and substantial performance gains across both IID and OOD evaluation scenarios against baseline methods.
@@
Quantitative Results Analysis.  Table 1 presents comprehensive performance comparisons across both ScienceWorld and ALFWorld environments. Our EPO enhancement demonstrates substantial improvements when integrated with existing RL methods. Notably, PPO+EPO achieves a remarkable 152.1% improvement in averaged success rates ($\overline{Succ.}$) on ScienceWorld IID tasks, significantly outperforming agent-specialized RL methods including GiGPO(53.4%) and RLVMR(67.2%). The dramatic improvement of PPO stems from addressing its inherent tendency toward aggressive policy updates that can cause severe entropy collapse in multi-turn interactions, particularly problematic in ScienceWorld’s sparse reward environment where maintaining exploration is crucial. Our entropy smoothing regularization directly mitigates this issue by stabilizing policy entropy across sequential turns. In contrast, GRPO’s more conservative update mechanism makes it less susceptible to such entropy instabilities, resulting in more modest but consistent improvements (9.5% on ALFWorld IID tasks) when combined with EPO. We emphasize the $\overline{Succ.}$ metric as it represents performance averaged across multiple evaluation episodes, providing a more robust and fair comparison by reducing variance from individual runs.
Training Dynamics Analysis.  Figure 2 illustrates the training dynamics and validation performance, revealing key insights into our method’s effectiveness. The training reward curves demonstrate that EPO-enhanced methods achieve substantially higher reward accumulation while maintaining superior stability. In ScienceWorld, PPO+EPO reaches approximately 2x higher training rewards (15 vs. 8) with smooth monotonic trajectories, while ALFWorld shows consistent improvements with GRPO+EPO maintaining steady upward trends throughout training. The validation performance curves reveal the most significant improvements in early training phases and overall convergence behavior. On ScienceWorld, our EPO variants achieve rapid convergence to high success rates (>0.8 for both IID and OOD) within 40 training steps, compared to baseline methods that struggle to exceed 0.4 even after 100 steps. On ALFWorld, while baseline methods exhibit substantial oscillations and instability, our EPO-enhanced approaches demonstrate more consistent performance with reduced variance, particularly evident in the OOD evaluation where baselines frequently drop below 0.2 while EPO variants maintain performance above 0.4. The key observable pattern across both environments is the elimination of the characteristic oscillations between premature convergence and over-exploration that plague standard RL methods. This stability translates to more reliable training outcomes and enhanced generalization capabilities, directly validating our entropy regularization framework’s effectiveness in addressing the exploration-exploitation dilemma in multi-turn LLM agent training.
5.3 Ablation Study
@@
Figure 3: Ablation studies on entropy regularization components. (a-c) ScienceWorld comparison of EPO versus EPO-Base without entropy smoothing, demonstrating that smoothing is essential for stable convergence in sparse reward settings. (d-f) ALFWorld comparison of EPO with dynamic $\beta_{k}$ versus EPO W/O DW using constant $\beta$, showing that adaptive weighting significantly accelerates early training progress.
Figure 3 presents ablations on two key components: the entropy smoothing regularizer and its dynamic weighting coefficient $\beta_{k}$. On ScienceWorld, removing the smoothing regularizer (EPO-Base) severely degrades performance. EPO-Base exhibits delayed convergence with minimal rewards until step 40 and plateaus at 0.5-0.6 success rate, while full EPO achieves meaningful learning by step 20 and reaches 0.8-1.0 success rate, demonstrating a 50-60% improvement. ScienceWorld’s sparse reward structure induces severe oscillations between exploration and exploitation, which the smoothing regularizer effectively stabilizes by maintaining entropy within historical bounds defined by $\kappa_{l}$ and $\kappa_{r}$. On ALFWorld, replacing the dynamic coefficient $\beta_{k}$ with a constant weight (EPO W/O DW) yields similar final performance around 0.7-0.8 success rate but results in slower convergence during the initial 40-60 episodes and increased training variance. The adaptive schedule in Equation 7, which progresses from $\beta_{start}$ through unity to $\beta_{end}$, automatically modulates regularization intensity across training phases, thereby improving convergence efficiency while preserving final performance quality. For more detailed ablation study, it can refer to  Appendix B.3.

### v2

@@ 4.1 Entropy Regularization
To address the exploration-exploitation cascade failure, we first adapt entropy regularization to capture the temporal structure of multi-turn interactions. Unlike traditional RL where entropy is computed per-step, we recognize that in multi-turn environments, decisions compound across subsequent turns through shared policy parameters. Therefore, we compute entropy across all turns within each trajectory and average over the batch of trajectories. The entropy-regularized policy loss is formulated as $L^{ER}(\theta)=L^{MT}(\theta)-\lambda L^{H}(\theta)$, where $\lambda$ is the entropy coefficient, and the entropy loss is averaged over the batch of trajectories:
@@ 4.2 Entropy Smoothing Regularizer
While trajectory-level entropy regularization captures multi-turn structure, it cannot prevent the severe entropy oscillations caused by shared parameters and sparse rewards. To stabilize entropy dynamics, we introduce an entropy smoothing mechanism that anchors current entropy to historical averages. We maintain an entropy history window $\mathcal{W}_{k}=\{\bar{H}_{0},\ldots,\bar{H}_{m},\ldots,\bar{H}_{k-1}\}$ for RL step $k$, storing the average entropy $\bar{H}_{m}$ across all trajectories at the token level for each previous RL step $m$. The historical entropy reference is computed as the mean $\bar{H}^{\mathcal{W}_{k}}=\frac{1}{k}\sum_{m=0}^{k-1}\bar{H}_{m}$. This historical anchoring dampens the entropy fluctuations that destabilize training under sparse reward conditions. We apply a token-wise penalty based on acceptable entropy ranges relative to this historical average:
$\mathcal{P}_{n,t,i}=\left[\kappa_{l}\bar{H}^{W_{k}}-H_{n,t,i}\right]_{+}+\left[H_{n,t,i}-\kappa_{r}\bar{H}^{W_{k}}\right]_{+}$ (3)
where $x]_{+}=\max(0,x)$ denotes the ReLU function, and $\kappa_{l}$, $\kappa_{r}$ define the acceptable entropy corridor. By bounding entropy within historical averages, we prevent the severe oscillations that arise from shared parameter coupling under sparse rewards. Aggregating these penalties across all tokens, turns, and trajectories yields the smoothing loss:
$L^{smooth}(\theta)=\frac{1}{B}\sum_{n=1}^{B}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{|\tau_{n,t}|}\sum_{i=1}^{|\tau_{n,t}|}\alpha\mathcal{P}_{n,t,i}$ (4)
@@
where $\alpha$ provides the penalty weight for tokens with entropy outside the acceptable range. The complete entropy-smoothed policy optimization loss is then defined as $L^{EPO}(\theta)=L^{MT}(\theta)-\lambda[L^{H}(\theta)-\beta_{k}L^{smooth}(\theta)]$, where the dynamic coefficient $\beta_{k}$ follows an exponential schedule that adapts smoothing strength across training:
$\beta_{k}=1+e^{-\gamma\frac{k}{k_{mid}}}$ (5)
This adaptive schedule stabilizes entropy dynamics throughout training: it begins with stronger smoothing to establish controlled exploration, gradually relaxes as the policy stabilizes around mid-training ($k_{mid}=\lfloor K/2\rfloor$), and maintains sufficient regularization in later phases to ensure smooth convergence. The parameter $\gamma$ controls the decay rate of this transition. Overall, the policy is updated according to the following objective:
$L^{EPO}(\theta)=L^{MT}(\theta)-\lambda[L^{H}(\theta)-\beta_{k}L^{\text{smooth}}(\theta)].$ (6)
where $\lambda$ is the entropy coefficient controlling the overall strength of entropy regularization, and $\beta_{k}$ is a dynamic coefficient that adapts the smoothing strength across training phases. By decoupling the entropy terms from the policy loss, this formulation provides direct gradient signals $\nabla_{\theta}L^{H}(\theta)$ to guide exploration while the smoothing regularizer $L^{\text{smooth}}(\theta)$ ensures stability by anchoring entropy within historical bounds. Algorithm 1 presents the full optimization procedure, incorporating these components.
Algorithm 1 Entropy-regularized Policy Optimization (EPO)
@@
1:Policy $\pi_{\theta}$, entropy coefficient $\lambda$, penalty weight $\alpha$, corridor bounds $\kappa_{l},\kappa_{r}$, decay rate $\gamma$, total steps $K$
2:Initialize $\theta_{0}$, entropy history $\mathcal{W}_{0}\leftarrow\emptyset$
3:for $k=0,1,\ldots,K-1$ do
4:  Collect trajectories $\mathcal{D}_{k}=\{\tau_{j}\}_{j=1}^{B}$ using $\pi_{\theta_{\text{old}}}$
5:  Compute advantages $\hat{A}_{t}$ and policy loss $L^{MT}(\theta)$
6:  Compute entropy loss $L^{H}(\theta)$ via Equation 2
7:  if $k>0$ then
8:   $\bar{H}^{\mathcal{W}_{k}}\leftarrow\frac{1}{k}\sum_{m=0}^{k-1}\bar{H}_{m}$ {Entropy mean}
9:   Compute loss $L^{\text{smooth}}(\theta)$ via Equation 4
10:  else
11:   $L^{\text{smooth}}(\theta)\leftarrow 0$
12:  end if
13:  $\beta_{k}\leftarrow 1+\exp(-\gamma\cdot k/k_{mid})$ where $k_{mid}=\lfloor K/2\rfloor$
14:  Update $\theta$ by minimizing Equation 6
15:  $\mathcal{W}_{k+1}\leftarrow\mathcal{W}_{k}\cup\{\bar{H}_{k}\}$ {Append current batch entropy}
16:end for
17:Optimized policy parameters $\theta_{K}$
@@
Table 1: For PPO and GRPO baselines with our EPO method: better performance indicates improvement over baseline, worse performance indicates degradation. Highlighted values represent the best performance among other baseline methods. $\Delta$ shows the relative improvement (%) when applying our method. Results for other baseline methods (ReAct, AgentGym, SFT, GiGPO, RLVMR) are sourced from the RLVMR (zhang2025rlvmr) paper. We ran our own implementations of PPO, GRPO, and EPO, tuning hyperparameters across multiple trials to obtain stable results.
@@
ReAct (react) GPT-4o 45.4 - 49.2 - GPT-4o 57.3 - 66.0 -
ReAct (react) DeepSeek-R1 22.2 - 31.4 - DeepSeek-R1 68.8 - 70.2 -
ReAct (react) Qwen2.5-7B 7.8 - 11.3 - Qwen2.5-7B 23.1 - 28.5 -
AgentGym (agentgym) LLaMa2-7B 46.9 - 33.6 - LLaMa2-7B 76.6 - 63.3 -
@@
GiGPO (gigpo) Qwen2.5-7B 53.4 - 35.2 - Qwen2.5-7B 89.5 - 90.2 -
RLVMR (zhang2025rlvmr) Qwen2.5-7B 67.2 - 43.0 - Qwen2.5-7B 91.4 - 91.8 -
PPO (ppo) Qwen2.5-7B 64.6 38.4 58.3 39.1 Qwen2.5-3B 95.8 72.3 87.5 70.9
@@
GRPO (grpo) Qwen2.5-7B 93.8 81.6 91.7 80.9 Qwen2.5-3B 87.5 63.3 83.3 63.5
@@
5 Experiments
We validate EPO on ScienceWorld and ALFWorld, two sparse-reward benchmarks requiring 30+ turn interactions. Our experiments examine: (1) performance gains over existing RL methods, (2) necessity of each proposed component via ablation, and (3) failure modes of alternative entropy control strategies. Results demonstrate that EPO effectively eliminates cascade failure, transforming unstable training into smooth convergence.
5.1 Experiments Setup
Benchmark. We evaluate on two challenging benchmarks that require different reasoning capabilities, ScienceWorld (sciworld) and ALFWorld (alfworld). ScienceWorld focuses on text-based scientific experimentation, demanding systematic hypothesis testing and structured exploration. ALFWorld is an embodied environment containing 4,639 household task instances across six categories, requiring multi-step decision-making and spatial reasoning. To improve the generalizability of our approach across multiple scenarios, we finetune the foundation model directly on the environment using RL, rather than employing trajectory finetuning for initialization.
Baselines. We conduct comprehensive comparisons across multiple paradigms: (1) Prompting-based approaches such as ReAct (react) that utilize in-context learning without parameter optimization; (2) Trajectory-based and platform methods including supervised fine-tuning (SFT) through expert trajectory imitation and AgentGym (agentgym) which provides a unified framework with behavioral cloning and self-evolution mechanisms; (3) General reinforcement learning approaches encompassing standard on-policy methods (PPO (ppo), GRPO (grpo)); (4) Agent RL approaches including recent methods specifically designed for agent training (GiGPO (gigpo), RLVMR (zhang2025rlvmr)). Our proposed EPO method is architected as a general enhancement framework that can be seamlessly integrated with existing RL paradigms, as exemplified through our PPO+EPO and GRPO+EPO implementations.
@@ 5.1 Experiments Setup
Figure 2: Training dynamics and generalization performance. (a-c) ScienceWorld results comparing PPO and PPO+EPO across training rewards, IID success rate, and OOD success rate. (d-f) ALFWorld results comparing GRPO and GRPO+EPO under identical metrics. EPO eliminates the characteristic entropy oscillations of cascade failure, transforming unstable training into smooth convergence with substantial gains in both IID and OOD scenarios.
Implementation Details. To optimize performance for each benchmark’s complexity, we employ Qwen2.5-3B-Instruct for ALFWorld and the larger Qwen2.5-7B-Instruct for ScienceWorld’s more complex scientific reasoning tasks. Constrained by computational resources, we adopt a single foundation model per task whose size is sufficient to ensure proper convergence, as smaller models consistently fail to converge. We conduct our own implementations and experiments for our proposed method, PPO, and GRPO baselines across three random seeds to ensure statistical reliability. Results for other baseline methods (ReAct, AgentGym, SFT, GiGPO, RLVMR) are sourced from the RLVMR’s (zhang2025rlvmr) paper. To account for their different convergence characteristics, we trained the model for 120 RL steps on ScienceWorld and 150 steps on ALFWorld.
5.2 Performance Comparison
@@
Quantitative Results Analysis. Table 1 presents comprehensive performance comparisons across both ScienceWorld and ALFWorld environments. Our EPO enhancement demonstrates substantial improvements when integrated with existing RL methods. Notably, PPO+EPO achieves a remarkable 152.1% improvement in averaged success rates ($\overline{Succ.}$) on ScienceWorld IID tasks, significantly outperforming agent-specialized methods including GiGPO(53.4%) and RLVMR(67.2%). This dramatic improvement stems from EPO’s ability to address the exploration-exploitation cascade failure: PPO’s aggressive policy updates amplify entropy oscillations under sparse rewards, which EPO’s smoothing regularizer effectively dampens by anchoring entropy within historical bounds. In contrast, GRPO’s group-relative advantage computation provides inherent stability against such oscillations, yielding more modest but consistent improvements (19.8% on ALFWorld IID) when combined with EPO. We emphasize the $\overline{Succ.}$ metric as it represents performance averaged across multiple evaluation episodes, providing a more robust comparison by reducing variance from individual runs.
[…]
@@
Figure 3: Ablation studies on entropy regularization components. (a-c) ScienceWorld comparison of EPO versus EPO-Base without entropy smoothing, demonstrating that smoothing is essential for stable convergence in sparse reward settings. (d-f) ALFWorld comparison of EPO with dynamic $\beta_{k}$ versus EPO W/O DW using constant $\beta$, showing that adaptive weighting accelerates early training progress.
(a) EPO-Base vs EPO-Decay
(b) Early vs Late Entropy
(c) EPO vs EA Success Rate
Figure 4: Model studies on ScienceWorld employing PPO with weighted entropy loss and entropy-based advantage shaping (cheng2025reasoning).
The key pattern across both environments is the elimination of severe entropy oscillations, the characteristic signature of cascade failure (Figure 1(c)), that destabilize standard RL methods. By maintaining entropy within historical bounds through our smoothing regularizer, EPO transforms previously unstable sparse-reward scenarios into smoothly converging optimization, validating our framework’s effectiveness in multi-turn LLM agent training.
5.3 Ablation Study
Figure 3 presents ablation studies on two key components: the entropy smoothing regularizer and its dynamic weighting coefficient $\beta_{k}$, evaluated across two base methods (PPO, GRPO) and two environments (ScienceWorld, ALFWorld).
Entropy Smoothing Regularizer. On ScienceWorld, removing the smoothing regularizer (EPO-Base) severely degrades performance. EPO-Base exhibits delayed convergence with minimal rewards until step 40 and plateaus at 0.5–0.6 success rate, while full EPO achieves meaningful learning by step 20 and reaches 0.8–1.0 success rate, demonstrating a 50–60% relative improvement. This gap directly reflects the cascade failure mechanism: ScienceWorld’s sparse rewards induce severe entropy oscillations that compound across turns, which the smoothing regularizer effectively dampens by anchoring entropy within the corridor $\kappa_{l},\kappa_{r}]$.
Dynamic Weighting Coefficient. On ALFWorld, replacing the adaptive $\beta_{k}$ with a constant weight (EPO W/O DW) yields comparable final performance (0.7–0.8 success rate) but results in slower convergence during the initial 40–60 episodes and increased training variance. The adaptive schedule (Equation 5) automatically modulates regularization intensity across training phases, applying stronger smoothing early to establish controlled exploration and relaxing mid-training as the policy stabilizes. This improves convergence efficiency without sacrificing final performance. See Appendix B.3 for detailed analysis.